OpenStudy (anonymous):

Hey what is the gcf of 24 and 196 @skateboarding

OpenStudy (anonymous):

4

OpenStudy (anonymous):

It is 4 use euclid's algorithm: http://codepad.org/AYWJBroy

OpenStudy (anonymous):

24=3*2*2*2; 196=4*7*7

OpenStudy (anonymous):

hi agdgdgdgwngo its me sk84 life